# Enhancing Metabolic Syndrome Prediction with Hybrid Data Balancing and Counterfactuals

### v1

@@ IV Discussion
Our research makes several important contributions to addressing these challenges. First, we developed a novel hybrid data balancing framework MetaBoost that combines SMOTE, ADASYN, and CTGAN and shows promise to perform just as well as other methods. It has the potential to perform better than other methods under different, more relevant circumstances which require a more nuanced, sophisticated approach. This approach demonstrated robust results with an accuracy of 86.88% and an F1 score of 86.52%, particularly noteworthy given the complexity of MetS prediction. Second, our probabilistic analysis revealed crucial insights into MetS risk factors, with elevated blood glucose showing the highest likelihood (85.5%) among MetS patients and triglycerides demonstrating the strongest predictive power (posterior probability 74.9%). Third, our counterfactual analysis provided actionable insights for clinical intervention, identifying triglycerides (46.7%) and blood glucose (50.3%) as the most influential features for shifting patients between risk categories.
@@ V Conclusion
This study systematically evaluated and optimized machine learning models for predicting Metabolic Syndrome (MetS), addressing critical challenges such as class imbalance, data scarcity, and methodological inconsistencies. By leveraging advanced data balancing techniques, including random oversampling (ROS), SMOTE, ADASYN, and CTGAN, as well as a novel hybrid framework combining these methods, we demonstrated its potential to be used in other settings where it has the potential to perform well and possibly better than traditional approaches. Our probabilistic analysis provided valuable insights into the relationships between key risk factors and MetS, revealing that elevated blood glucose had the highest likelihood (85.5%) among MetS patients, while triglycerides exhibited the strongest predictive power with a posterior probability of 74.9%. The counterfactual analysis further enhanced interpretability, identifying triglycerides (46.7%) and blood glucose (50.3%) as the most frequently altered features in counterfactual examples.
The hybrid data balancing framework (MetaBoost), combining SMOTE, ADASYN, and CTGAN, achieved an accuracy of 86.88% and an F1 score of 86.52%. This approach not only shows potential for model generalizability but also provides a unique, scalable solution for handling imbalanced datasets in healthcare analytics. Additionally, the decision boundary visualizations and feature-specific change rates derived from counterfactual analysis can offer actionable insights for clinicians.
In conclusion, this study advances the methodological rigor of MetS prediction by integrating advanced ML techniques with probabilistic and counterfactual analyses. The findings underscore the potential of hybrid data balancing and interpretable ML models in addressing the global public health burden of metabolic syndrome. Future work should focus on validating these approaches in diverse populations and exploring real-time deployment in clinical settings to enable early intervention and personalized care.

### v2

@@ IV Discussion
Our research makes several important contributions to addressing these challenges. First, we developed a novel hybrid data balancing framework, MetaBoost, that chooses and combines a subset of SMOTE, ADASYN, and CTGAN and shows promise to perform better than other methods. This approach demonstrated robust results with an accuracy of 87.1% and an F1 score of 0.868 when ADASYN and CTGAN are chosen, particularly noteworthy given the complexity of MetS prediction.
@@ V Conclusion
This study systematically evaluated and optimized machine learning models for predicting Metabolic Syndrome (MetS), addressing critical challenges such as class imbalance, data scarcity, and methodological inconsistencies. By leveraging advanced data balancing techniques, including random oversampling (ROS), SMOTE, ADASYN, and CTGAN, as well as a novel hybrid framework combining these methods, we demonstrated its potential to be used in other settings where it has the potential to perform well and possibly better than traditional approaches. The counterfactual analysis enhanced the interpretability of the solution, identifying triglycerides and blood glucose as the most frequently altered features in counterfactual examples.
The hybrid data balancing framework (MetaBoost), combining ADASYN and CTGAN, achieved an accuracy of 87.1% and an F1 score of 0.868. This approach not only shows potential for model generalizability but also provides a unique, scalable solution for handling imbalanced datasets in healthcare analytics. Additionally, the counterfactual analysis can offer actionable insights for clinicians.
In conclusion, this study advances the methodological rigor of MetS prediction by integrating advanced ML techniques with probabilistic and counterfactual analyses. The findings underscore the potential of hybrid data balancing and interpretable ML models in addressing the global public health burden of metabolic syndrome.